How many different​ 10-letter words​ (real or​ imaginary) can be formed from the following​ letters?
w,o,e,k,t,u,o,q,f,v
1
Expert's answer
2019-10-21T10:31:35-0400

We can take one of 10 letters for first letter of word, then we can take one of 9 remaining letters for second letter of word, and so on.

We obtain "10\cdot 9\cdot\ldots\cdot 1=10!" combinations of letters.

Since we have only two equal letters o, every combination we count twice. So we need to "10!" divide by 2, and we obtain every combination, which is counted only one time.

Answer: "\frac{10!}{2}"